OpenStudy (anonymous):

if K(-7,-4) and P(5,2) find the slope of the line to KP 21 1 1/2 2

OpenStudy (anonymous):

It's C) 1/2

OpenStudy (anonymous):

\[m = \frac{ 2 - (-4) }{ 5 - (-7) } = \frac{ 6 }{12 } = \frac{ 1 }{ 2 }\]